OpenStudy (ledah):

simplify 9x^3+6x^2+3x divided by 3x

OpenStudy (ledah):

3x^4+6x^3+x2 3x^2-2xy-2 3x^2+2x+1 None of the above

OpenStudy (maheshmeghwal9):

\[\frac{9x^3+6x^2+3x}{3x}.\]\[\implies \frac{9x^3}{3x}+\frac{6x^2}{3x}+\frac{3x}{3x}.\]

OpenStudy (moongazer):

divide 9x^3+6x^2+3x by 3x one by one per addend.

OpenStudy (maheshmeghwal9):

\[\implies 3x^2+2x+1\]

OpenStudy (ledah):

Okay I get it now.

OpenStudy (ledah):

thanks a bunch

OpenStudy (maheshmeghwal9):

yw:)